# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

### v2

@@ 4.2 The Overall Algorithm via Certificates and Caching Them
Now, we prove that this algorithm outputs a $(1-\varepsilon)$-approximation matching.
Lemma 11.
Given graph $G$ with $n$ vertices, Section 4.2 maintains an $(1-\varepsilon$)-approximate maximum matching for $G$ in the dynamic model.
Proof.
In this algorithm, we maintain $O(\log^{2}n)$ contraction of graph $G$, where $G$ is the updated graph in the current batch of $t$ updates. If the maximum matching size of $G$ is $\mu$, by Proposition 6, there exists a contraction $c$ with $n_{c}$ vertices that preserves the maximum matching with $1-\delta$ factors and $\mu=O(n_{c})$. Denote $\delta$ a parameter sufficiently smaller than $\varepsilon$. We call $c$ a preserving contraction of $G$. This means that by Lemma 10, Section 4.2 outputs a $(1-\delta)$-approximation matching for $G^{\prime}$, that is $(1-\varepsilon)$-approximation matching for $G$. Note that, by running Proposition 4 on the maintained reduction, we find the preserving contraction of $G$.This concludes the $(1-\varepsilon)$-approximate. ∎
4.3 Running Time Analysis
Now, we explore the time per update for Section 4.2.
Lemma 12.
If we set the threshold of Section 4.2 equal to $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}$, then the time spent on single contraction $G^{\prime}$ in both Section 4.2 is $O(\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)})$ per each update, where $\mu$ is the maximum matching size in $G^{\prime}$.
Proof.
Fix a contracted graph $G^{\prime}=(V^{\prime},E^{\prime})$ with $n$ vertices. We explore the algorithm running time for this graph.
First, let us determine the time the algorithm spends finding the size of the certificate matching, $\mu$, after $\varepsilon^{2}n$ updates. Note that we run Proposition 4 after each $\varepsilon^{2}n$ vertex update, this algorithm outputs $\mu$, which is a constant approximation of the maximum matching size of the graph. Moreover, this algorithm certifies the existence of a matching of size $\mu$ in the graph, the running time for which is $\widetilde{O}(n\operatorname{poly}(1/\varepsilon))$.
By Proposition 3, after each $\varepsilon^{2}n$ update we spend $\widetilde{O}_{\varepsilon}(n)$ on finding the appropriate subsets of vertices such as $U$, including matching of size at least $\varepsilon n$.
Now, we calculate the time Section 4.2 takes going through the set $S_{1}$ and $S_{2}$, after $\varepsilon^{2}n$ updates to find a matching $M^{\prime}$. Note that, after $\varepsilon^{2}n$ updates we go over the set $S_{1}\cup S_{2}$ to find an active matching. We add an active edge to the matching if none of the endpoints was added before it. The algorithm outputs a maximal matching this way. The total running time per $\varepsilon^{2}n$ updates is $|S_{1}|+|S_{2}|$.
Finally, we explore the running time of Section 4.1. Let us assume that at the end of a phase of Section 4.2 we have $S_{2}=\{M_{1}\cup\dots\cup M_{l}\}$, We prove in Lemma 10 that these matchings are edge-disjoint. Note that the total edge-disjoint matchings in $G^{\prime}$ can be more than $l$ matching, $M_{1},\dots,M_{l}$ are all the matchings we find during the $t$ edge updates. We assume that the total edge-disjoint matchings in $G^{\prime}$ is $l_{t}$. We label these matchings as $M_{1},\dots,M_{l},M_{l+1},\dots,M_{l_{t}}$. Let us assume that for $i\in\{1,\dots,l\}$ we denote $d_{i}$ the average degree of matching $M_{i}$ in the edge set of $\{M_{i}\cup\dots\cup M_{l_{t}}\}$. Denoting by $M^{\prime}$ a maximal matching in $S_{1}\cup S_{2}$ found in Section 4.2, we prove that if Section 4.1 in graph $G^{\prime}\setminus V(M^{\prime})$, takes $O(n^{2+\varepsilon}/d)$ to output matching $M_{i}$ then we have $O(d)=d_{i}$. Note that by Lemma 8 we know that the average degree of matching $M_{i}$ in $G^{\prime}\setminus(S_{1}\cup S_{2})$ is $O(d)$. Note that, for matching $M_{1},\dots,M_{i-1}$, either their edges are removed from the graph or we deactivate the vertices that made a large matching. This implies that no edges from $M_{1},\dots,M_{i-1}$ remain in the graph. If we assumed that an edge $e$ was still in the graph with both endpoints not deactivated, we could add this edge to $M^{\prime}$, this is in contradiction with the fact that $M^{\prime}$ is a maximal matching. This implies that $d_{i}=O(d)$. This results in the conclusion that the running time of Section 4.1 is less than $n^{2+\varepsilon}/d_{i}$.
This definition of matching average degree motivates us to define ORS graphs. Recall that the ORS graph is a set of matching $M_{1},\dots M_{l}$ that each matching is induced in the set of matching with higher indices. We then upper bound the running time of our algorithm by the size of ORS graphs according to Lemma 13.
Now, let us compute the total running time per update. Note that since we add a matching of size $\varepsilon\mu$ to $S_{2}$ after $\varepsilon^{2}n$ updates, $|S_{2}|<|S_{1}|$. Let us assume that we add $M_{1},\dots M_{l}$ to $S_{2}$ by running Section 4.1. The total running time per update is:
$\frac{\frac{|S_{1}|}{\varepsilon^{2}n}\cdot(|S_{1}|+|S_{2}|+\widetilde{O}(n))+n^{2+\varepsilon}\cdot\sum_{i=1}^{l}1/d_{i}}{t}.$
Using Lemma 13 we have the running time as:
$\frac{\frac{|S_{1}|}{\varepsilon^{2}n}\cdot(|S_{1}|+|S_{2}|+\widetilde{O}(n))+n^{2+\varepsilon}\cdot O(\mathsf{ORS}_{n}(\varepsilon\mu))\log n}{t}.$
Now, if we set the threshold as $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}$ we have:
$|S_{2}|,|S_{1}|<t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)},$
This is because we could add a matching of size $\varepsilon\mu$ to $S_{2}$ after $\varepsilon^{2}n$ updates, and the total insertions are less than the total updates. The total running time per update is:
$O(\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}).\qed$
Lemma 13.
Given a set of matching $M_{1},\dots,M_{t}$ with distinct edges, of size at least $r$ in graph $G$, for any index $i\in[1,t]$ let us define $d_{i}$ as the average degree of $V(M_{i})$ in $M_{1}\cup\dots\cup M_{i}$. For any $\alpha\in(0,1)$, we have $\sum_{i=1}^{t}1/d_{i}=O(\mathsf{ORS}_{n}(r(1-\alpha))\cdot\log n)$.
Proof.
Note that the maximum average degree of a matching $M$ is $n-1$, and the minimum is 1. Therefore, if we partition the input matchings into $\log n=k$ groups, the average degree of any matching in each partition is at most within 2 factors of the minimum average degree.
We first prove that if we have $t^{\prime}$ matchings $M^{\prime}_{1},\dots,M^{\prime}_{t^{\prime}}$ a subset of $\{M_{1},\dots,M_{t}\}$, and their average degree as $d^{\prime}_{1}\star,\dots,d^{\prime}_{t^{\prime}}$ that for any $i\in[1,t^{\prime}]$, $1/d^{\prime}_{i}$ is in a $2$ factor of $\min_{i=1}^{t^{\prime}}1/d^{\prime}_{i}$. For any $\alpha\in(0,1)$, we have $\sum_{i=1}^{t^{\prime}}1/d^{\prime}_{i}=O(\mathsf{ORS}_{n}(r(1-\alpha)))$.
To prove this, we first sample some of the matchings and the edges that are not induced in proper subsets of matchings. We prove there are plenty number of matchings remaining that are induced within their corresponding set of matchings.
Let us fix a small variable $\delta<\varepsilon,\delta,\varepsilon\in(0,1)$. We set $H=M^{\prime}_{1}\cup...\cup M^{\prime}_{t^{\prime}}$.
  • Step 1. Let us define $x_{e}$ for any edge $e=(u,v)$ appearing in matching $M^{\prime}_{\alpha_{j}}$ as follows
    $\sum_{i\in\{j,j+1,\dots,t^{\prime}\};u,v\in V^{2}(M_{\alpha_{i}})}\frac{1}{d^{\prime}_{\alpha_{i}}}.$
    We define $E_{1}=\{e|x_{e}>1/\varepsilon\}$ We remove any matching from $H$ with more than $2\varepsilon r$ edges in $E_{1}$ Let us call the remaining set of matchings $H_{1}$.
  • Step 2. Let us pick a matching $M^{\prime}_{i}\in H$ with probability $\frac{\delta}{2d^{\prime}_{i}}$. Let us assume that at this step, we pick the following $k$ matchings $M^{\prime}_{\alpha_{1}},\dots,M^{\prime}_{\alpha_{k}}$. we define this subset of the graph as $H_{2}$.
  • Step 3. Now, we prone any matching $M^{\prime}_{i}$ chosen in $H_{1}$ and has less than $(1-\delta)cn$ edges induced in $(M^{\prime}_{1}\cup\dots\cup M^{\prime}_{i})\cap(H_{2})$. Let us call the remaining graph $H_{3}$.
To see that the claim is true, first, we need to prove that in Step 1 we remove at most $\varepsilon$ fraction of the edges in $H$. Second, we prove that $H_{2}$ is large enough.
Let us assume that after Step 1 we remove more than $\varepsilon rt$ of the edges in $H$ from the total $rt^{\prime}$ edges. Therefore, we have
$\sum_{e\in E_{1}}x_{e}>rt^{\prime}.$
To compute $\sum_{e\in H}x_{e}$, we notice how much a matching $M^{\prime}_{\alpha_{i}}$ contributes to the total sum. Note that since the average degree of $M^{\prime}_{\alpha_{i}}$ in $\{M^{\prime}_{\alpha_{1}},\dots,M^{\prime}_{\alpha_{i}}\}$ is at most $d^{\prime}_{\alpha_{i}}$ therefore, the total contribution of matching $M^{\prime}_{\alpha_{i}}$ is $r\cdot d^{\prime}_{\alpha_{i}}\cdot\frac{1}{d^{\prime}_{\alpha_{i}}}$. In conclusion, the total sum of $x_{e}$ for any edge $e\in H$ is at most
$\sum_{i\in\{1,\dots,t^{\prime}\}}d^{\prime}_{\alpha_{i}}\cdot r/d^{\prime}_{\alpha_{i}}=rt^{\prime},$
which is a contradiction suggesting that we remove at most $\varepsilon rt^{\prime}$ edges.
Now that we only removed at most $\varepsilon$ fraction of the edges, note that $H_{1}$ contains at least half of the matchings in $H$. This is because if more than $|H|/2$ matchings had more than $2\varepsilon r$ edges removed, then the total number of edges in $E_{1}$ would be more than $\varepsilon rt^{\prime}$ which is a contradiction.
We aim to prove that $H_{3}$ contains $\delta|H_{1}|/2$ matchings. Let us fix a matching $M\in H_{1}$. Note that for any remaining edge $e$ in $M$, we have $x_{e}<1/\varepsilon$. Therefore, the probability that $e$ belongs to another matching is less than $\delta/2\varepsilon$ since we picked each matching $M^{\prime}_{i}$ with probability $\delta/2d^{\prime}_{i}$. This implies that the expected number of edges in $E(^{\prime})\cap E(H_{3})$ is more than $(1-2\varepsilon)(1-\delta/2\varepsilon)r$. Now, by applying Markov’s inequality to the random variable $r(1-2\varepsilon)-|E(M)\cap E(H_{3})|$, here we assume that the maximum number of edges is $r(1-2\varepsilon)$, and we get the following.
$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]\leq\frac{\operatorname{\textbf{E}}[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|M\in H_{2}]}{(1-2\varepsilon)(\delta/\varepsilon)r}.$
Since $\operatorname{\textbf{E}}[r(1-2\varepsilon)-|E(M)\cap E(H_{3})||M\in H_{2}]\leq((1-2\varepsilon)\delta/2\varepsilon)r$ we have
$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]\leq\frac{(\delta/2\varepsilon)r}{(\delta/\varepsilon)r}=1/2.$
Therefore, we have
$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\leq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]>1/2,$
which means
$\Pr[|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(1-\delta/\varepsilon)r|M\in H_{2}]>1/2.$
Multiplying with the probability that $M\in H_{2}$, we get that the probability of matching to have more than $(1-2\varepsilon)(1-\delta/\varepsilon)r$ edges and be selected is more than $\delta/(4\varepsilon d^{\prime}_{i})$.
Now, summing over all matchings in $H_{1}$ we get
$\operatorname{\textbf{E}}[|H_{3}|]\geq(\delta/(4\varepsilon))\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}.$
Since $H_{3}$ is an ORS graph we get
$(4\varepsilon)\mathsf{ORS}_{n}((1-2\varepsilon)(1-\delta/\varepsilon)r)/\delta\geq\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}.$
Now, using that for any $j\in\{1,..,t^{\prime}\}$ we have $1/d^{\prime}_{j}\leq 2\min_{i=1}^{t^{\prime}}1/d^{\prime}_{i}$, for any $i,j\in\{1,...,t^{\prime}\}$ we get $1/d^{\prime}_{i}\leq 2/d^{\prime}_{j}$.Therefore,
$\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}\geq\left(\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}\right)/2.$
Now, we give an upper bound for $\sum_{i,M_{i}\in H}1/d^{\prime}_{i}=\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}+\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}$. We have the following.
$(12\varepsilon)\mathsf{ORS}_{n}((1-2\varepsilon)(1-\delta/\varepsilon)r)/\delta\geq 3\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}\geq\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}+\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}.$
Setting $\delta=\varepsilon/2,\varepsilon=4\alpha/13$ we have:
$\sum_{i,M^{\prime}_{i}\in H}1/d^{\prime}_{i}=O(\mathsf{ORS}_{n}((1-\alpha)r)).$
Note that we showed an upper bound for the sum of the inverse average degree of each matching in one partition. Summing up the inverse average degree for all the partitions we get:
$\sum_{i=1}^{t}1/d_{i}=O(\mathsf{ORS}_{n}((1-\alpha)r)\cdot\log n.\qed$
Lemma 14.
If we set the threshold of Section 4.2 equal to $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n)}$, then the total time per update is $O(\operatorname{poly}\log n)\cdot\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}$.
Proof.
We run the $O(\log^{2}n)$ parallel version of Section 4.2, and maintaining the contracted graphs takes $\operatorname{poly}\log n$ time. We will maintain the number of parallel edges in $G^{\prime}$ between any two contracted vertex in $\operatorname{poly}\log n$ time. Denote the set of contractions as $\mathcal{C}$ in one phase of the algorithm.
Note that if the size of a contraction $c$ is $n_{c}$, and the maximum matching in $G$ is of order $n_{c}$ we run the algorithm for this contraction. Using Proposition 6 there is a reduction that preserves that matching in $G$ with a factor of $1-\varepsilon$. Now, integrating over the contractions in $\mathcal{C}$ we find the corresponding contraction by running Proposition 4. If the maximum matching in contraction $c$ is $M_{c}$, this algorithm certifies the existence of a matching of size $|M_{c}|/2<\mu\leq|M_{c}|$ in at least one contraction with $n_{c}$ vertices. On the other hand, Proposition 3 outputs subsets with a matching of size at least $\varepsilon n_{c}$. This means that the $\varepsilon$-approximation matching found on each subset determined by Proposition 3 is at least of size $\varepsilon^{2}n$.
Now, since we run Section 4.2 on all contractions, nevertheless they contain the maximum matching or not, assuming $n_{c}$ is the number of vertices in $c$, we have the total running time as:
$\sum_{c\in\mathcal{C}}\sqrt{n_{c}^{1+\varepsilon}\log n_{c}\cdot\mathsf{ORS}_{n_{c}}(\Theta(\varepsilon^{2}n_{c}))}\leq O(\operatorname{poly}\log n)\cdot\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}.\qed$
Now, using Lemma 11 and Lemma 14 the proof of Theorem 1 is complete.
@@ 5.1 Linear Matchings
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 16, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
@@
$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 2.)
@@
Instead of the expected value of $Y_{M}$, we need some (weak) concentration. Take $\varepsilon=\delta c^{2}/2$ and note from 3 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$. We have
@@
We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 4 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
@@
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 16). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎
@@
$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 6.)
@@
$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 5.)

### v3

@@ 4.4 Runtime Analysis of Theorem 1
Note that for each matching $M_{i}\in H_{3}$, we have $|M_{i}|\geq(1-\varepsilon)(1-2\varepsilon)r\geq(1-3\varepsilon)r$ by construction. Additionally, $M_{i}$ is an induced matching in $\bigcup_{j<i;M_{j}\in H_{3}}M_{j}$, again, by construction. Therefore, $H_{3}$ is an ORS graph with matchings of size $(1-3\varepsilon)r$. This implies that
$\mathsf{ORS}_{n}((1-3\varepsilon)r)\geq E[|H_{3}|]\geq\Omega(\varepsilon^{2}t/d).$
Moving the terms and recalling that $\varepsilon=\alpha/3$, we get that
$t=O\left(\frac{d}{\alpha^{2}}\cdot\mathsf{ORS}_{n}((1-\alpha)r)\right).\qed$
Now, we are ready to prove Lemma 11.
Proof of Lemma 11.
Given $M_{1},\dots M_{t}$, we partition the matchings to $\log n$ groups $P_{1},\dots,P_{\log n}$ where $M_{j}$ is in the $i$-th group if $2^{i-1}\leq d_{j}<2^{i}$. Define $d^{\prime}_{j}:=|E[P_{i}]\cap V^{2}(M_{j})|/|M_{j}|$. Note that by definition for any matching $M_{j}\in P_{i}$ we have $d^{\prime}_{j}\leq d_{j}\leq 2^{i}$. Therefore, by Claim 12 we have
$\sum_{i\in[\log n]}\sum_{j:M_{j}\in P_{i}}1/d_{j}$ $\leq\sum_{i\in[\log n]}|P_{i}|/2^{i-1}$
$\leq O\left(\frac{1}{\alpha^{2}}\mathsf{ORS}_{n}((1-\alpha)r)\right)\cdot\sum_{i\in[\log n]}2^{i}/2^{i-1}$ (Applying Claim 12 with $d=2^{i}$, we get $|P_{i}|\leq O\left(\frac{1}{\alpha^{2}}\mathsf{ORS}_{n}((1-\alpha)r)2^{i}\right)$ which is plugged here.)
$=O\left(\frac{1}{\alpha^{2}}\mathsf{ORS}_{n}((1-\alpha)r)\cdot\log n\right).$
The proof is thus complete. ∎
@@ 5.1 Linear Matchings
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
@@
$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
@@
Instead of the expected value of $Y_{M}$, we need some (weak) concentration. Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$. We have
@@
We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
@@
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎
@@
$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
@@
$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)